OpenStudy (anonymous):

Which is the equation to the line passing through points (2, -7) and (5, -1)? y = 2x - 8 y = - 2x + 9 y = 2x - 11 y = - 2x + 6

OpenStudy (haseeb96):

first take out the the slope by this formula m= y2 - y1 /x2 -x1 let (x1,y1)=(2,-7) (x2,y2)=(5,-1)

OpenStudy (haseeb96):

so the slope is m=2 alright!!!

OpenStudy (haseeb96):

now use the point slope formula here pick the slope and only any one point from the two given points (y-y1)=m(x-x1) now put them in it and get the answer

OpenStudy (haseeb96):

let (x1,y1)=(5,-1) m=2 then y+1=2(x-5) y+1 =2x-10 y=2x-10 -1 y=2x-11 is the correct answer

OpenStudy (haseeb96):

@kellyconnelly have u got it????

OpenStudy (anonymous):

yes thanks you were helpful:)

OpenStudy (haseeb96):

welcome my dear @kellyconnelly

OpenStudy (anonymous):

What is the equation of a vertical line passing through (1, 7)? y = 7 x = 1 y = 1 x = 7

OpenStudy (anonymous):

for that question would you do the same thing?

OpenStudy (anonymous):

it only gives me one point so idk how to find slope?

OpenStudy (haseeb96):

here it is not given either the vertical line is on the y - axis or far from it

OpenStudy (anonymous):

soooo what do i do??

OpenStudy (haseeb96):

x=1 is the correct answer

OpenStudy (driftracer305):

@kellyconnelly did u guys figure it out?

OpenStudy (haseeb96):

A vertical line passing through the point (4, -3) means that for any value of y, x is always going to be equal to 4; the formula may be generally written as x = a where a is the x-intercept.. An x-intercept is the value of x when y = 0. As we already stated, for any value of y, x = 4. Therefore, this vertical line will pass through the point (4,0). You can write the equation as x = 4.

OpenStudy (haseeb96):

. Graphing this equation in a graphing calculator requires you to "trick" the graphing calculator because in the "graph" menu you only have the option of writing in equations in the form y =, which you can't do for this problem

OpenStudy (anonymous):

wut

OpenStudy (driftracer305):

lol........ it is so simple

OpenStudy (anonymous):

show me how you would solve it please

OpenStudy (driftracer305):

\[-7 = 2(2) - 11.......................-7 = 4 -11\] \[-1 = 2(5) -11.................... -1 = 10-11 \]

OpenStudy (driftracer305):

the answer would be C @kellyconnelly ..........

OpenStudy (haseeb96):

@kellyconnelly @driftracer305 is correct because here it is asking about vertical line and we know that on thhe vertical line the x=0 then y=1 clapping for @driftracer305

OpenStudy (anonymous):

so its y=1 not x=1 ????

OpenStudy (haseeb96):

yes it is y=1 my apologize i forget the major thing that it is asking about the vertical line

OpenStudy (anonymous):

oh its fine thank you!!